A typical pool is 3.0 m deep. What is the water pressure at the bottom of the pool?

1
Expert's answer
2022-01-06T09:06:59-0500
"P=\rho gh=1000\ \dfrac{kg}{m^3}\times 9.8\ \dfrac{m}{s^2}\times 3\ m=29400\ Pa."